OpenStudy (anonymous):

An opinion poll asks an SRS of 1500 adults, "Do you happen to jog?" Suppose that the population proportion who jog (a parameter) is p = 0.15. To estimate p, we use the proportion p-hat in the sample who answer "Yes." Justify the use of a normal approximation and find the following probabilities. a) P(p > 0.16) b) P(0.14 < p < 0.16)

OpenStudy (lucylu252):

a) P(p > 0.16) means find the probability that the sample proportion is greater than 0.16 b) P(0.14 < p < 0.16) means find the probability that the sample proportion lies between 0.14 and 0.16 Use z distribution for answering the question z = (O - E) / SEp O denotes observed or sample proportion or p-hat E denotes expected or population proportion or p p = 0.15 and n = 1500 SEp denotes standard error of the sample proportion = sqrt (p*(1-p)/n) = sqrt (0.15*0.85/1500) = 0.0092 a) z = (0.16 - 0.15) / 0.0092 = + 1.09 Required probability = P(p > 0.16) = P(z > 1.09) The area under the standard normal curve right to the z value indicates the required probability. This area lies in the extreme right tail of the normal curve. = 0.5000 (total area on the right side of p) - 0.3621(area between p and z) = 0.1379 b) z1 = (0.14 - 0.15) / 0.0092 = - 1.09 z2 = (0.16 - 0.15) / 0.0092 = + 1.09 Required probability = P(0.14 < p < 0.16) = P(- 1.09 < z < 1.09) The area under the standard normal curve between the two z values indicate the required probability. = 0.3621 (area corresponding to z1 which lies on the left side of p) + 0.3621 (area corresponding to z2 which lies on the right side of p) = 0.7242